OpenStudy (anonymous):

A block is at rest on a plank whose angle can be varied. The angle is gradually increased from 0 deg. At 28.4°, the block starts to slide down the incline, traveling 2.80 m down the incline in 1.50 s. Calculate the coefficient of static friction between the block and the plank. Calculate the kinetic coefficent of friction between the block and the plank.

OpenStudy (anonymous):

I already have the first part....I cannot figure out the kinetic coefficent part....please help!

OpenStudy (perl):

what is the formula for coefficient of static friction

OpenStudy (anonymous):

you need to figure out 'a' ( resultant acceleration down the incline )from the data traveling 2.80 m down the incline in 1.50 s and if the coefficient of kinetic friction be x then we can have g sin 28.4 - x g cos 28.4 = a

OpenStudy (anonymous):

so 9.8sin28.4=4.66 and 9.8cos28.4=8.62

OpenStudy (anonymous):

4.66-28.4x=a?

OpenStudy (anonymous):

is x = 1.125?

OpenStudy (anonymous):

opps I mean 4.66-8.62*1.125=a

OpenStudy (anonymous):

first find a

OpenStudy (anonymous):

im not sure how

OpenStudy (anonymous):

soooo a is equal to 1.244?

OpenStudy (anonymous):

I need to visualize this